### v1

# 22footnotetext: To whom correspondence should be addressed

Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors. After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity. This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available in the appendix where the $n=6,\dots,9$ ICT decomposition matrices are obtained in <0.1s, 0.5s, 1s, 3s, 11s, and 4m32s, respectively.
Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant). Spherical tensors serve as the starting point for EGNN based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
$T^{(n;j;q)}=H^{(n;j;q)}T^{(n)},$ (1)
where $T^{(n;j;q)}$ is an ICT, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, $j$ is called the weight of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$). The decomposition matrices $H$ should be constructed such that $T^{(n;j;q)}$ are irreducible. The challenge lies in the fact that $H$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible). Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2). Therefore, the matrices for high-order orthogonal ICT decomposition are of great importance. Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
$\rho(g)f(T)=f(\rho(g)T)$ (2)
where $g\in G$, and $\rho$ is the matrix representation. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
In this work, one of our contributions is to efficiently obtain ICT decomposition matrices for $6\leq n\leq 9$. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverted change of basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector. Thus, we can multiply it by a diagonalized matrix with only zeros and ones on the diagonals. This operation eliminates the other spherical vectors and retains only the desired one. The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space. The product of these three matrices forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs. This approach takes less than one second on CPUs to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$. We also show that this process can be further accelerated. Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that a specific decomposition matrix can be generated via the product of the path matrix and its transpose.
Method Ranks Orthogonality Time/Space complexity
coope1965irreducible 3 No calculated by hand
andrews1982irreducible 4 No calculated by hand
dincckal2013orthonormal 3 Yes calculated by hand
bonvicini2024irreducible 5 No factorial
Ours 6-9 Yes exponential
Table 1: Comparison with previous works.
Another challenge this work addresses is the construction of bases for the equivariant spaces. With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight). These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal. Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space. We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces. Additionally, we demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities. This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
This work has three main contributions:
  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$.
  • We characterize the bases of equivariant spaces and propose a method to efficiently obtain these bases, overcoming the redundancy in pearce2023brauer’s method.
  • We extend our approach to efficiently obtain bases for equivariant spaces operating between different spaces.
These contributions enable EGNNs to process high-rank ICTs and flexibly design equivariant layers between desired spaces. Furthermore, this work benefits the physics and chemistry communities by facilitating theoretical analysis based on high-rank ICTs.
The following contents are organized as follows. In Section 2, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact. We introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The code and visualizations are shown in the appendix.
2 Preliminaries
2.1 Representations and group equivariance
We first recall some basic facts. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. We first focus on the $O(3)$ group (3D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion. Consider a vector $v\in V$. The rotation and reflection operations should be able to act on it. Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation. It holds that
$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
for any $g_{1}$, $g_{2}\in G$. The rotation $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in V$. In this sense, this place is also called a $G$-invariant space or subspace. To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal. It is important to note that the co-domain of $\rho$ can differ, meaning it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a function $f:X\to Y$ is equivariant if and only if
$\rho_{Y}f(x)=f(\rho_{X}(x)).$
We only consider $f$ as a linear function, and thus it is represented as a matrix. The set of all such $f$ constitutes a vector space. We sometimes denote this space as ${\rm End}_{O(3)}(X)$ if $X=Y$, and as ${\rm Hom}_{O(3)}(X,Y)$ otherwise. Meanwhile, we say that the function $f$ is invariant if $\rho_{Y}=1$. Since $X$ is not necessarily the same as $Y$, this actually makes finding the function $f$ relatively straightforward. Specifically, to fulfill
$\rho_{Y}(g)M=M\rho_{X}(g),$
we can simply let
$\rho_{Y}(g)=M\rho_{X}(g)M^{-1}.$
However, this approach is very time-consuming in the design of EGNNs. The matrix $M$ here is learnable in the networks. During training, we must repeatedly compute the inverse of $M$, which is computationally expensive. Even if we could afford this, the results of the $M_{i}v_{i}$ is not addable in general, as the co-domains of different $M_{i}$ vary. Moreover, we cannot guarantee that $M$ is actually invertible without adding restriction on the determinants.
Considering that linear combinations involve a tremendous number of inversion operations, this process becomes intractable. Therefore, we first focus on the simpler case where $X=Y$. We will later generalize this to the case where $X\neq Y$, but $Y$ is known and remains consistent across different $M$. When it is necessary to emphasize the change in space, we refer to a linear mapping $M$ as $G$-$(X,Y)$-equivariant.
2.2 Irreducible representations and spherical spaces
We then introduce the definition of reducible and irreducible representation, in a matrix view:
Definition 1
A group homomorphism $\rho\in{\rm Hom}_{G}(\mathbbm{R}^{k})$ is said to be reducible, if there is an invertible matrix $P$, such that for every $g\in G$, we have
$P\rho(g)P^{-1}={\rm diag}\big{(}\rho_{1}(g),\rho_{2}(g)\big{)}$
for some non-trivial $\rho_{1}$ and $\rho_{2}$.
The $3\times 3$ rotation and reflection matrices are irreducible. More generally,

### v2

@@
 We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry.
-The Python code is available in the appendix where the $n=6,\dots,9$ ICT decomposition matrices are obtained in <0.1s, 0.5s, 1s, 3s, 11s, and 4m32s, respectively.
+The Python code is available in https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s, respectively.
 Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
@@
 One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations.
-Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant).
+Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$.
 Spherical tensors serve as the starting point for EGNN based on irreducible representations.
@@
 The ICTs are obtained through the ICT decomposition matrices $H$,
-$T^{(n;j;q)}=H^{(n;j;q)}T^{(n)},$ (1)
-where $T^{(n;j;q)}$ is an ICT, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, $j$ is called the weight of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$).
-The decomposition matrices $H$ should be constructed such that $T^{(n;j;q)}$ are irreducible.
+${\rm vec}\left(T^{(n;j;q)}\right)=H^{(n;j;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
+where $T^{(n;j;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $j$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$).
+The decomposition matrices $H^{(n;j;q)}$ should be constructed such that $T^{(n;j;q)}$ are irreducible.
+A formal definition is introduced in Section LABEL:sec:ictdecomp.
 The challenge lies in the fact that $H$ of rank $n>5$ are unknown.
@@
 Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
-$\rho(g)f(T)=f(\rho(g)T)$ (2)
+$\rho(g)f\left({\rm vec}(T)\right)=f\left(\rho(g){\rm vec}(T)\right)$ (2)
 where $g\in G$, and $\rho$ is the matrix representation.
@@
 Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that a specific decomposition matrix can be generated via the product of the path matrix and its transpose.
-Method Ranks Orthogonality Time/Space complexity
+Method Rank Orthogonality Time/Space complexity
 coope1965irreducible 3 No calculated by hand
@@
 Ours 6-9 Yes exponential
-Table 1: Comparison with previous works.
+Table 1: Comparison with previous works. The complexity is discussed in terms of rank. The complexity of this work will be polynomial if we focus on the dimensions.
 Another challenge this work addresses is the construction of bases for the equivariant spaces.
@@
 This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
-This work has three main contributions:
-  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$.
+This work has four main contributions:
+  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, compared to predecessors, as shown in Table 1.
+  • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
   • We characterize the bases of equivariant spaces and propose a method to efficiently obtain these bases, overcoming the redundancy in pearce2023brauer’s method.
@@
 The acknowledgments are presented after the main text.
-The code and visualizations are shown in the appendix.
+The visualizations are shown in the appendix.
+A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure 1.
+Table 2: Mathematical symbols and their meanings
+Symbol Meaning Symbol Meaning
+$\lVert\cdot\rVert_{col}$ The $L_{2}$-norm of an arbitrary column of $\cdot$ $G$ An arbitrary group
+$SO(n)$ Special orthogonal group on $n$-dimensional space $SU(n)$ Special unitary group on $n$-dimensional space
+$O(n)$ Orthogonal group on $n$-dimensional space $U(n)$ Unitary group on $n$-dimensional space
+$\mathcal{S}$ A vector space $\rho_{\mathcal{S}}$ A group representation
+$\mathcal{S}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{S}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
+$T$ An arbitrary tensor $G\to\mathcal{S}$ $T^{(n)}$ A Cartesian tensor of rank $n$
+$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{S}}$ The set of all possible paths to decompose a tensor in space $\mathcal{S}$
+$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{S};j;q)}$ A general ICT decomposition matrix of space $\mathcal{S}$, weight $j$ and index $q$
+$C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
+$C^{l_{1}l_{2}l_{o}}$ A CG tensor with input weights $l_{1}$, $l_{2}$, and output weight $l_{o}$ $\hat{C}^{l_{1}l_{2}l_{o}}$ A CG matrix by merging the first two dimensions of the original CG tensor
+$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
+$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A representation weight $l$ and parity $p$ irreducible representation acts on
+${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
+irr Returns all weights of irreducible representations in a decomposition ${\rm dim}$ The number of dimensions of a space
+$(i_{1}\overset{k_{1}}{\to}i_{2}$ $\overset{k_{2}}{\to}\dots)$ A path where $k_{*}$ are bridge numbers (omitted if all being $1$) $(l,p)$ An irreducible representation space of weight $l$ and parity $p$
+$\otimes$ Tensor product $\oplus$ Direct sum
+$\odot$ Contraction ##\## Number of elements in a set
 2 Preliminaries
-2.1 Representations and group equivariance
-We first recall some basic facts.
-A thorough introduction to group theory can be found in inui2012group and brocker2013representations.
-We first focus on the $O(3)$ group (3D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion.
-Consider a vector $v\in V$.
-The rotation and reflection operations should be able to act on it.
-Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation.
-It holds that
-$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
-for any $g_{1}$, $g_{2}\in G$.
-The rotation $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$.
-Sometimes, one may also refer to the space of $v$ as a representation space.
-If the results of the group action remain in the same space, i.e., $\rho(g)v\in V$.
-In this sense, this place is also called a $G$-invariant space or subspace.
-To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal.
-It is important to note that the co-domain of $\rho$ can differ, meaning it can also be represented by matrices of other dimensions.
-In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$.
-Now, it suffices to define equivariance.
-We claim that a function $f:X\to Y$ is equivariant if and only if
-$\rho_{Y}f(x)=f(\rho_{X}(x)).$
-We only consider $f$ as a linear function, and thus it is represented as a matrix.
-The set of all such $f$ constitutes a vector space.
-We sometimes denote this space as ${\rm End}_{O(3)}(X)$ if $X=Y$, and as ${\rm Hom}_{O(3)}(X,Y)$ otherwise.
-Meanwhile, we say that the function $f$ is invariant if $\rho_{Y}=1$.
-Since $X$ is not necessarily the same as $Y$, this actually makes finding the function $f$ relatively straightforward.
-Specifically, to fulfill
-$\rho_{Y}(g)M=M\rho_{X}(g),$
-we can simply let
-$\rho_{Y}(g)=M\rho_{X}(g)M^{-1}.$
-However, this approach is very time-consuming in the design of EGNNs.
-The matrix $M$ here is learnable in the networks.
-During training, we must repeatedly compute the inverse of $M$, which is computationally expensive.
-Even if we could afford this, the results of the $M_{i}v_{i}$ is not addable in general, as the co-domains of different $M_{i}$ vary.
-Moreover, we cannot guarantee that $M$ is actually invertible without adding restriction on the determinants.
-Considering that linear combinations involve a tremendous number of inversion operations, this process becomes intractable.
-Therefore, we first focus on the simpler case where $X=Y$.
-We will later generalize this to the case where $X\neq Y$, but $Y$ is known and remains consistent across different $M$.
-When it is necessary to emphasize the change in space, we refer to a linear mapping $M$ as $G$-$(X,Y)$-equivariant.
-2.2 Irreducible representations and spherical spaces
-We then introduce the definition of reducible and irreducible representation, in a matrix view:
-Definition 1
-A group homomorphism $\rho\in{\rm Hom}_{G}(\mathbbm{R}^{k})$ is said to be reducible, if there is an invertible matrix $P$, such that for every $g\in G$, we have
-$P\rho(g)P^{-1}={\rm diag}\big{(}\rho_{1}(g),\rho_{2}(g)\big{)}$
-for some non-trivial $\rho_{1}$ and $\rho_{2}$.
-The $3\times 3$ rotation and reflection matrices are irreducible. More generally,
+2.1 Clebsch-Gordan coefficients
+A review on irreducible representations and spherical spaces is in the appendix.
+The term weights in this paper refer to the order of irreducible representations.
+An important tool utilized in our work is the Clebsch-Gordan coefficients (CG coefficients).
+They are real numbers, and describe how to couple vectors from two spherical spaces into a new one.
+There are some rules for this coupling.
+First, given two vectors in weights $l_{1}$ and $l_{2}$ spherical spaces, the weights of the resulting vectors can only lie in $|l_{1}-l_{2}|\leq l_{o}\leq l_{1}+l_{2}$.
+For each index $m_{o}$ of the obtained vector $v_{l_{o}}\in(l=l_{o})$, it is calculated by homogeneous quadratic polynomials.
+Specifically, we have
+$v^{(l_{o})}_{m_{o}}=\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1})}v_{m_{2}}^{(l_{2})},$ (3)
+and if we consider O(n) instead of SO(n), parity $p$ should also be considered.
+The parity is a property additional to the rank and weight, being $1$ or $-1$.
+Having $-1$ parity means that the representation will be inverse when the space is inverse, and vice versa.
+For example, the natural representations of O(n) always have $p=-1$.
+We will introduce the case that $p=1$ in the next section.
+Taking the parity into consideration, we can guarantee O(n) equivariance, and equation (3) becomes
+$v^{(l_{o};p_{o})}_{m_{o}}=\mathbbm{1}_{(p_{o}=p_{1}p_{2})}\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1};p_{1})}v_{m_{2}}^{(l_{2};p_{2})},$
+where $\mathbbm{1}_{(exp)}$ is the indicator function, outputting $1$ if $exp$ holds true, or 00 otherwise.
+Our discussion will be the same no matter we consider $O(n)$ or $SO(n)$.
+But since Section LABEL:sec:extension, we need to take care of the parity which will play some roles in the further discussion.
+The CG coefficients are highly sparse, since $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}=0$ if $m_{o}\neq m_{1}+m_{2}$.
+The coefficients are identical to Wigner-3j symbols up to multiplicative factors, so we will adopt Wigner-3j for the implementation.
+The evaluation of CG coefficients is nearly $\mathcal{O}(1)$ (if we do not consider the few calculations for the closed-form formula, which is cheap enough).
+To obtain the coefficients of $O(3)/SO(3)$, one often leverages the relation between $SU(2)$ and $SO(3)$: $SU(2)$ is a double cover of $SO(3)$, so the CG matrices of $SU(2)$ and $SO(3)$ are same up to a change-of-basis, from complex to real.
+The closed-form formulae for $SU(2)$ CG coefficients are already known.
+Thus, this process can be efficiently evaluated without obvious overhead.
+Figure 1: The mind map of our method. The blocks in orange color are aligned with the three contributions of this work.
+2.2 Tensors and tensor product spaces
+To formally define a tensor, we need to consider vector space $V$ and covector space $V^{*}$.
+A tensor product is a bilinear map $\otimes:\mathcal{S}\times\mathcal{S}^{\prime}\to\mathcal{S}\otimes\mathcal{S}^{\prime}$.
+The basis of $\mathcal{S}\otimes\mathcal{S}^{\prime}$ is the set of $\hat{x}_{i}\otimes\hat{y}_{j}\in\mathcal{S}\otimes\mathcal{S}^{\prime}$, where $\hat{x}_{i}$ and $\hat{y}_{j}$ are basis vectors of $\mathcal{S}$ and $\mathcal{S}^{\prime}$, respectively.
+That is why one implements outer product as tensor product, as it meets the definition of tensor product.
+A tensor $T:V^{\otimes r}\otimes V^{*\otimes s}\to F$, where $F$ is some field, is a mathematical object that converts things between different tensor product spaces.
+For example, the representation $\rho(g)$ is a $(r=1,s=1)$-tensor that receives a vector $v$ which is a $(r=1,s=0)$-tensor, and output another $(r=1,s=0)$-tensor $\rho(g)v$, as $r$ and $s$ vanish correspondingly.
+Tensor product can help us build up the ranks of tensors, e.g., $V\otimes V\otimes V$ is a $(r=3,s=0)$ space.
+The representation space is accordingly $(r=3,s=3)$.
+Conversely, contraction can reduce the ranks of tensors.
+We express these operations via some index notations.
+For example, $T_{i_{1}i_{2}}=T^{\prime}_{i_{1}}T^{\prime\prime}_{i_{2}}$ is a tensor product, and $T_{i_{1}}=\sum_{j_{1}}T^{\prime}_{i_{1}j_{1}}T^{\prime\prime}_{j_{1}}$ is a contraction.
+Obviously, such operations can be mixed together, and we often omit the index of the output tensors and agree $i_{k}$ will be placed at the $k$th index of the output tensor by default.
+We also agree that $j_{1}$ and $j_{2}$ will vanish due to the contraction.
+We will not use Einstein summation convention in this work to keep the conventions of machine learning community, and to avoid ambiguous during derivations.
+One should note that contractions can be expressed in term of matrix representation.
+Consider $T=\sum_{j_{1},j_{2}}T^{\prime}_{i_{1}j_{1}j_{2}}T^{\prime\prime}_{j_{1}j_{2}}$, we have
+$\begin{bmatrix}T_{1}\\ T_{2}\end{bmatrix}=\begin{bmatrix}T^{\prime}_{111}&T^{\prime}_{112}&T^{\prime}_{121}&T^{\prime}_{122}\\ T^{\prime}_{211}&T^{\prime}_{212}&T^{\prime}_{221}&T^{\prime}_{222}\end{bmatrix}\begin{bmatrix}T^{\prime\prime}_{11}\\ T^{\prime\prime}_{12}\\ T^{\prime\prime}_{21}\\ T^{\prime\prime}_{22}\end{bmatrix}.$
+It is useful when we need to obtain the inverse of some linear operations. Also, contractions and tensor products are both $O(n)$-equivariant.
+2.3 CG matrices and the parentage scheme
+Figure 2: The parentage scheme. The numbers in the circle represent multiplicity. The sum of each column equals equation (LABEL:eq:multiplicity).
+The $O(3)$ group has its representations in tensor product space $(\mathbbm{R}^{3})^{\otimes n}$, which are some $3^{n}\times 3^{n}$ matrices, or tensor product of $n$ $(r=1,s=1)$-tensors, which depends on one’s view.
+If we interpret them as matrices, then the objects they act on are simply $3^{n}$-dimensional vectors.
+The tensor product spaces $(\mathbbm{R}^{3})^{\otimes n}$ for $n>1$ are generally reducible.
+Recall CG coefficients, two irreducible spherical vectors $v$ and $v^{\prime}$ coupling can also be viewed as that they first form a rank 2 tensor, and then decompose into different new spherical vectors.
+These operations can also be viewed as a product of matrices and vectors.
+In this context, the CG coefficients $\hat{C}^{l_{1}l_{2}l_{o}}$ form a $(2l_{1}+1)(2l_{2}+1)\times(2l_{o}+1)$ matrix, which we call a CG matrix.
+If we reshape it to a $(2l_{1}+1)\times(2l_{2}+1)\times(2l_{o}+1)$ tensor, then we instead call it an $(l_{1},l_{2},l_{o})$-tensor $C^{l_{1}l_{2}l_{o}}$.
+Here are some important properties,
+Lemma 1 (Orthogonality)
+Given an $(l_{1},l_{2},l_{o})$-CG tensor $C^{l_{1}l_{2}l_{o}}$, we have
+$\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k^{\prime}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}kj_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}k^{\prime}j_{2}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{kj_{1}j_{2}}C^{l_{1}l_{2}l_{o}}_{k^{\prime}j_{1}j_{2}}=0$
+for $k\neq k^{\prime}$. More generally, different CG tensors with the same $l_{1}$ and $l_{2}$, $l_{1}$ and $l_{3}$, or $l_{2}$ and $l_{3}$ are also orthogonal in this way.

### v3

@@
 The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors.
-After decades of research, we recently achieve an explicit ICT decomposition for $n=5$ (bonvicini2024irreducible) with factorial time/space complexity.
+After decades of research, bonvicini2024irreducible recently achieves an explicit ICT decomposition for $n=5$ with factorial time/space complexity.
 This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices.
@@
 We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry.
-The Python code is available in https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s, respectively.
+The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
 Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
@@
 However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive.
-In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs), making them more computationally friendly for most deep learning frameworks.
+In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks.
 The ICTs are obtained through the ICT decomposition matrices $H$,
-${\rm vec}\left(T^{(n;j;q)}\right)=H^{(n;j;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
-where $T^{(n;j;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $j$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $j$ is sometimes greater than $1$).
-The decomposition matrices $H^{(n;j;q)}$ should be constructed such that $T^{(n;j;q)}$ are irreducible.
+${\rm vec}\left(\hat{T}^{(n;l;q)}\right)=H^{(n;l;q)}{\rm vec}\left(T^{(n)}\right),$ (1)
+where $\hat{T}^{(n;l;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $l$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $l$ is sometimes greater than $1$).
+The decomposition matrices $H^{(n;l;q)}$ should be constructed such that $\hat{T}^{(n;l;q)}$ are irreducible.
 A formal definition is introduced in Section LABEL:sec:ictdecomp.
@@
 The decomposition matrices for $n=4$ have been established by andrews1982irreducible.
-However, only very recently have we obtained the decomposition matrices for $n=5$ (bonvicini2024irreducible).
-Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weight equals their rank (zaverkin2024higher).
+However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$.
+Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
 It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2).
@@
 $\rho(g)f\left({\rm vec}(T)\right)=f\left(\rho(g){\rm vec}(T)\right)$ (2)
-where $g\in G$, and $\rho$ is the matrix representation.
+where $g\in G$, $T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$, and $\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the matrix representation.
 This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$.
@@
 Thus, it is also essential to identify a basis to reduce the number of parameters required.
-In this work, one of our contributions is to efficiently obtain ICT decomposition matrices for $6\leq n\leq 9$.
+In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$.
 Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors.
@@
 Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis.
-In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverted change of basis can bring us back to Cartesian space.
+In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverse change-of-basis can bring us back to Cartesian space.
 The challenge lies in finding such a change-of-basis matrix.
@@
 Thus, we can multiply it by a diagonalized matrix with only zeros and ones on the diagonals.
-This operation eliminates the other spherical vectors and retains only the desired one.
+This operation eliminates the other spherical vectors and retains only the desired ones.
 The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space.
-The product of these three matrices forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs.
-This approach takes less than one second on CPUs to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$.
+The product of these three matrices—the change-of-basis matrix, the diagonalized matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs.
+This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz.
+to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$.
 We also show that this process can be further accelerated.
-Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that a specific decomposition matrix can be generated via the product of the path matrix and its transpose.
+Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that decomposition matrices can be generated via the product of the path matrices and their transpose, with no inverse operation involved.
 Method Rank Orthogonality Time/Space complexity
@@
 Ours 6-9 Yes exponential
-Table 1: Comparison with previous works. The complexity is discussed in terms of rank. The complexity of this work will be polynomial if we focus on the dimensions.
+Table 1: Comparison with previous works.
+The complexity is discussed in terms of rank.
+The complexity of this work will be polynomial if we focus on the dimensions of the tensor product spaces.
+In contrast, the work by bonvicini2024irreducible has a complexity of $\mathcal{O}(\log(d)!)$, which is slower than polynomial.
 Another challenge this work addresses is the construction of bases for the equivariant spaces.
@@
 We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces.
-Additionally, we demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities.
+We demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities.
 This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
-This work has four main contributions:
-  • We propose a method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, compared to predecessors, as shown in Table 1.
+In summary, this work has four main contributions:
+  • We propose a non-numerical method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
   • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
-  • We characterize the bases of equivariant spaces and propose a method to efficiently obtain these bases, overcoming the redundancy in pearce2023brauer’s method.
-  • We extend our approach to efficiently obtain bases for equivariant spaces operating between different spaces.
+  • We propose to efficiently obtain orthogonal bases of equivariant spaces, whereas the method proposed by pearce2023brauer yields a spanning set, requiring an expensive post-processing step to extract a basis that is not orthogonal in general.
+  • We extend our approach to efficiently obtain bases of equivariant spaces mapping between different spaces, including spherical tensor product spaces.
 These contributions enable EGNNs to process high-rank ICTs and flexibly design equivariant layers between desired spaces.
@@
 In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices.
-In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$.
-We provide concluding remarks in Section LABEL:sec:conclusion, and discuss the potential social impact in Section LABEL:sec:boardersocialimpact.
-We introduce the reproducibility details in Section LABEL:sec:repro.
-The acknowledgments are presented after the main text.
-The visualizations are shown in the appendix.
-A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure 1.
+In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well
 Table 2: Mathematical symbols and their meanings
@@
 $P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
-$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A representation weight $l$ and parity $p$ irreducible representation acts on
+$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A vector on which weight $l$ and parity $p$ irreducible representation acts
 ${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
@@
 $\odot$ Contraction ##\## Number of elements in a set
+as the theory for $SO(n)$, $O(n)$, and $SU(n)$.
+We provide concluding remarks in Section LABEL:sec:conclusion, and introduce the reproducibility details in Section LABEL:sec:repro.
+The acknowledgments are presented after the main text.
+The visualizations are shown in the appendix.
+A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure LABEL:fig:mindmap.
 2 Preliminaries
-2.1 Clebsch-Gordan coefficients
-A review on irreducible representations and spherical spaces is in the appendix.
-The term weights in this paper refer to the order of irreducible representations.
-An important tool utilized in our work is the Clebsch-Gordan coefficients (CG coefficients).
-They are real numbers, and describe how to couple vectors from two spherical spaces into a new one.
-There are some rules for this coupling.
-First, given two vectors in weights $l_{1}$ and $l_{2}$ spherical spaces, the weights of the resulting vectors can only lie in $|l_{1}-l_{2}|\leq l_{o}\leq l_{1}+l_{2}$.
-For each index $m_{o}$ of the obtained vector $v_{l_{o}}\in(l=l_{o})$, it is calculated by homogeneous quadratic polynomials.
-Specifically, we have
-$v^{(l_{o})}_{m_{o}}=\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1})}v_{m_{2}}^{(l_{2})},$ (3)
-and if we consider O(n) instead of SO(n), parity $p$ should also be considered.
-The parity is a property additional to the rank and weight, being $1$ or $-1$.
-Having $-1$ parity means that the representation will be inverse when the space is inverse, and vice versa.
-For example, the natural representations of O(n) always have $p=-1$.
-We will introduce the case that $p=1$ in the next section.
-Taking the parity into consideration, we can guarantee O(n) equivariance, and equation (3) becomes
-$v^{(l_{o};p_{o})}_{m_{o}}=\mathbbm{1}_{(p_{o}=p_{1}p_{2})}\sum_{m_{o}=m_{1}+m_{2}}C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}v_{m_{1}}^{(l_{1};p_{1})}v_{m_{2}}^{(l_{2};p_{2})},$
-where $\mathbbm{1}_{(exp)}$ is the indicator function, outputting $1$ if $exp$ holds true, or 00 otherwise.
-Our discussion will be the same no matter we consider $O(n)$ or $SO(n)$.
-But since Section LABEL:sec:extension, we need to take care of the parity which will play some roles in the further discussion.
-The CG coefficients are highly sparse, since $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}=0$ if $m_{o}\neq m_{1}+m_{2}$.
-The coefficients are identical to Wigner-3j symbols up to multiplicative factors, so we will adopt Wigner-3j for the implementation.
-The evaluation of CG coefficients is nearly $\mathcal{O}(1)$ (if we do not consider the few calculations for the closed-form formula, which is cheap enough).
-To obtain the coefficients of $O(3)/SO(3)$, one often leverages the relation between $SU(2)$ and $SO(3)$: $SU(2)$ is a double cover of $SO(3)$, so the CG matrices of $SU(2)$ and $SO(3)$ are same up to a change-of-basis, from complex to real.
-The closed-form formulae for $SU(2)$ CG coefficients are already known.
-Thus, this process can be efficiently evaluated without obvious overhead.
-Figure 1: The mind map of our method. The blocks in orange color are aligned with the three contributions of this work.
-2.2 Tensors and tensor product spaces
-To formally define a tensor, we need to consider vector space $V$ and covector space $V^{*}$.
-A tensor product is a bilinear map $\otimes:\mathcal{S}\times\mathcal{S}^{\prime}\to\mathcal{S}\otimes\mathcal{S}^{\prime}$.
-The basis of $\mathcal{S}\otimes\mathcal{S}^{\prime}$ is the set of $\hat{x}_{i}\otimes\hat{y}_{j}\in\mathcal{S}\otimes\mathcal{S}^{\prime}$, where $\hat{x}_{i}$ and $\hat{y}_{j}$ are basis vectors of $\mathcal{S}$ and $\mathcal{S}^{\prime}$, respectively.
-That is why one implements outer product as tensor product, as it meets the definition of tensor product.
-A tensor $T:V^{\otimes r}\otimes V^{*\otimes s}\to F$, where $F$ is some field, is a mathematical object that converts things between different tensor product spaces.
-For example, the representation $\rho(g)$ is a $(r=1,s=1)$-tensor that receives a vector $v$ which is a $(r=1,s=0)$-tensor, and output another $(r=1,s=0)$-tensor $\rho(g)v$, as $r$ and $s$ vanish correspondingly.
-Tensor product can help us build up the ranks of tensors, e.g., $V\otimes V\otimes V$ is a $(r=3,s=0)$ space.
-The representation space is accordingly $(r=3,s=3)$.
-Conversely, contraction can reduce the ranks of tensors.
-We express these operations via some index notations.
-For example, $T_{i_{1}i_{2}}=T^{\prime}_{i_{1}}T^{\prime\prime}_{i_{2}}$ is a tensor product, and $T_{i_{1}}=\sum_{j_{1}}T^{\prime}_{i_{1}j_{1}}T^{\prime\prime}_{j_{1}}$ is a contraction.
-Obviously, such operations can be mixed together, and we often omit the index of the output tensors and agree $i_{k}$ will be placed at the $k$th index of the output tensor by default.
-We also agree that $j_{1}$ and $j_{2}$ will vanish due to the contraction.
-We will not use Einstein summation convention in this work to keep the conventions of machine learning community, and to avoid ambiguous during derivations.
-One should note that contractions can be expressed in term of matrix representation.
-Consider $T=\sum_{j_{1},j_{2}}T^{\prime}_{i_{1}j_{1}j_{2}}T^{\prime\prime}_{j_{1}j_{2}}$, we have
-$\begin{bmatrix}T_{1}\\ T_{2}\end{bmatrix}=\begin{bmatrix}T^{\prime}_{111}&T^{\prime}_{112}&T^{\prime}_{121}&T^{\prime}_{122}\\ T^{\prime}_{211}&T^{\prime}_{212}&T^{\prime}_{221}&T^{\prime}_{222}\end{bmatrix}\begin{bmatrix}T^{\prime\prime}_{11}\\ T^{\prime\prime}_{12}\\ T^{\prime\prime}_{21}\\ T^{\prime\prime}_{22}\end{bmatrix}.$
-It is useful when we need to obtain the inverse of some linear operations. Also, contractions and tensor products are both $O(n)$-equivariant.
-2.3 CG matrices and the parentage scheme
-Figure 2: The parentage scheme. The numbers in the circle represent multiplicity. The sum of each column equals equation (LABEL:eq:multiplicity).
-The $O(3)$ group has its representations in tensor product space $(\mathbbm{R}^{3})^{\otimes n}$, which are some $3^{n}\times 3^{n}$ matrices, or tensor product of $n$ $(r=1,s=1)$-tensors, which depends on one’s view.
-If we interpret them as matrices, then the objects they act on are simply $3^{n}$-dimensional vectors.
-The tensor product spaces $(\mathbbm{R}^{3})^{\otimes n}$ for $n>1$ are generally reducible.
-Recall CG coefficients, two irreducible spherical vectors $v$ and $v^{\prime}$ coupling can also be viewed as that they first form a rank 2 tensor, and then decompose into different new spherical vectors.
-These operations can also be viewed as a product of matrices and vectors.
-In this context, the CG coefficients $\hat{C}^{l_{1}l_{2}l_{o}}$ form a $(2l_{1}+1)(2l_{2}+1)\times(2l_{o}+1)$ matrix, which we call a CG matrix.
-If we reshape it to a $(2l_{1}+1)\times(2l_{2}+1)\times(2l_{o}+1)$ tensor, then we instead call it an $(l_{1},l_{2},l_{o})$-tensor $C^{l_{1}l_{2}l_{o}}$.
-Here are some important properties,
-Lemma 1 (Orthogonality)
-Given an $(l_{1},l_{2},l_{o})$-CG tensor $C^{l_{1}l_{2}l_{o}}$, we have
-$\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k}C^{l_{1}l_{2}l_{o}}_{j_{1}j_{2}k^{\prime}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}kj_{2}}C^{l_{1}l_{2}l_{o}}_{j_{1}k^{\prime}j_{2}}=\sum_{j_{1},j_{2}}C^{l_{1}l_{2}l_{o}}_{kj_{1}j_{2}}C^{l_{1}l_{2}l_{o}}_{k^{\prime}j_{1}j_{2}}=0$
-for $k\neq k^{\prime}$. More generally, different CG tensors with the same $l_{1}$ and $l_{2}$, $l_{1}$ and $l_{3}$, or $l_{2}$ and $l_{3}$ are also orthogonal in this way.
+2.1 Irreducible representations and spherical spaces
+We first recall some basic facts.
+A thorough introduction to group theory can be found in inui2012group and brocker2013representations.
+We first focus on the $O(3)$ group (3-D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion.
+Consider a vector $v\in\mathcal{S}$.
+The rotation and reflection operations should be able to act on it.
+Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation.
+It holds that
+$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
+for any $g_{1}$, $g_{2}\in G$, and * is the group multiplication.
+The group element $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$.
+Sometimes, one may also refer to the space of $v$ as a representation space.
+If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{S}$.
+In this sense, this place is also called a $G$-invariant space or subspace.
+To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal.
+It is important to note that the co-domain of $\rho$ can differ, meaning that it can also be represented by matrices of other dimensions.
+In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$.
+Now, it suffices to define equivariance.
+We claim that a linear function $f:\mathcal{S}\to\mathcal{S}^{\prime}$ is equivariant if and only if
+$\rho_{\mathcal{S}^{\prime}}f(x)=f(\rho_{\mathcal{S}}(x)).$
+We first focus on the simpler case where $\mathcal{S}=\mathcal{S}^{\prime}$.
+We will later generalize this to the case where $\mathcal{S}\neq\mathcal{S}^{\prime}$.
+When it is necessary to emphasize the change in space, we refer to a linear mapping $f$ as $G$-$(\mathcal{S},\mathcal{S}^{\prime})$-equivariant.
+We then introduce the definition of reducible and irreducible representation.
+Definition 1
+A group representation $\rho_{\mathcal{S}}:G\to GL(\mathcal{S})$ is said to be reducible, if there is a non-zero subspace $\mathcal{S}^{\prime}\subset\mathcal{S}$, for which we have a representation $\rho_{\mathcal{S}^{\prime}}$, such that for every $g\in G$ and $v\in\mathcal{S}^{\prime}$, it holds that
+$\rho_{\mathcal{S}^{\prime}}(g)v\in\mathcal{S}^{\prime},$
+where we also call $\mathcal{S}^{\prime}$ a $G$-invariant subspace.
+Conversely, if there is no such subspace $\mathcal{S}^{\prime}$ and the corresponding representation on it for a given space $\mathcal{S}$, then we claim that it is irreducible.
+If a representation is irreducible, then the vectors space it acts on is also said to be irreducible. The $3\times 3$ rotation and reflection matrices are irreducible. More generally,

### v6

@@
 Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge.
-The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-order tensors.
+The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-rank tensors.
 After decades of research, bonvicini2024irreducible recently achieves an explicit ICT decomposition for $n=5$ with factorial time/space complexity.
-This work, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices.
-The path matrices are obtained via performing chain-like contraction with Clebsch-Gordan matrices following the parentage scheme.
+In this work we, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices.
+The path matrices are obtained via performing chain-like contractions with Clebsch-Gordan matrices following the parentage scheme.
 We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces.
 Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique.
+To the best of our knowledge, this is also the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
 We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry.
-The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
-Keywords: irreducible Cartesian tensor, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural networks, spherical harmonics
+The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
+Keywords: irreducible Cartesian tensor decomposition, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural network, equivariant linear mapping
 1 Introduction
@@
 Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$.
-Spherical tensors serve as the starting point for EGNN based on irreducible representations.
+Spherical tensors serve as the starting point for EGNNs based on irreducible representations.
 However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive.
-In contrast, ICTs are more efficient, as high-order representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks.
+In contrast, ICTs are more efficient, as high-rank representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks.
 The ICTs are obtained through the ICT decomposition matrices $H$,
@@
 A formal definition is introduced in Section LABEL:sec:ictdecomp.
-The challenge lies in the fact that $H$ of rank $n>5$ are unknown.
+The challenge lies in the fact that $H^{(n;l;q)}$ of rank $n>5$ are unknown.
 The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities.
@@
 However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$.
-Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-order ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
+Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-rank ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
 It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2).
-Therefore, the matrices for high-order orthogonal ICT decomposition are of great importance.
+Therefore, the matrices for high-rank orthogonal ICT decomposition are of great importance.
 Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
@@
 Thus, it is also essential to identify a basis to reduce the number of parameters required.
-In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$.
+In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting.
 Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors.
@@
 If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector.
-Thus, we can multiply it by a diagonalized matrix with only zeros and ones on the diagonals.
+Thus, we can multiply it by a diagonal matrix with only zeros and ones on the diagonal.
 This operation eliminates the other spherical vectors and retains only the desired ones.
 The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space.
-The product of these three matrices—the change-of-basis matrix, the diagonalized matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs.
-This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz.
-to compute cases for $n=6$, which is already faster than the current decomposition for $n=5$.
+The product of these three matrices—the change-of-basis matrix, the diagonal matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs.
+This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz to compute cases for $n=6$, which is already much faster than the current decomposition for $n=5$ by bonvicini2024irreducible.
 We also show that this process can be further accelerated.
 Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that decomposition matrices can be generated via the product of the path matrices and their transpose, with no inverse operation involved.
-Method Rank Orthogonality Time/Space complexity
-coope1965irreducible 3 No calculated by hand
-andrews1982irreducible 4 No calculated by hand
-dincckal2013orthonormal 3 Yes calculated by hand
-bonvicini2024irreducible 5 No factorial
-Ours 6-9 Yes exponential
-Table 1: Comparison with previous works.
-The complexity is discussed in terms of rank.
-The complexity of this work will be polynomial if we focus on the dimensions of the tensor product spaces.
-In contrast, the work by bonvicini2024irreducible has a complexity of $\mathcal{O}(\log(d)!)$, which is slower than polynomial.
+Method Rank Orthogonality Time and Space Complexity
+w.r.t. rank w.r.t. dimension
+coope1965irreducible $\leq 3$ No calculated by hand
+andrews1982irreducible $\leq 4$ No calculated by hand
+dincckal2013orthonormal $\leq 3$ Yes calculated by hand
+bonvicini2024irreducible $\leq 5$ No
+factorial  / sub-exponential
+Ours $\leq 9$ Yes
+exponential  / polynomial
+Table 1: Comparison with previous works. The time and space complexities are with respect to rank and dimension, respectively.
 Another challenge this work addresses is the construction of bases for the equivariant spaces.
@@
 This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
-In summary, this work has four main contributions:
-  • We propose a non-numerical method to efficiently calculate high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
+In summary, this work has five main contributions:
+  • We propose a non-numerical method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
   • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
-  • We propose to efficiently obtain orthogonal bases of equivariant spaces, whereas the method proposed by pearce2023brauer yields a spanning set, requiring an expensive post-processing step to extract a basis that is not orthogonal in general.
+  • We propose to directly obtain orthogonal bases of equivariant spaces, whereas the method proposed by pearce2023brauer yields a spanning set, requiring an expensive post-processing step to extract a basis that is not orthogonal in general.
+  • To the best of our knowledge, this is the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
   • We extend our approach to efficiently obtain bases of equivariant spaces mapping between different spaces, including spherical tensor product spaces.
@@
 Furthermore, this work benefits the physics and chemistry communities by facilitating theoretical analysis based on high-rank ICTs.
-The following contents are organized as follows.
-In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme.
-In Section LABEL:sec:problem, we formulate our problems: obtaining high-order ICT decomposition matrices and bases for arbitrary tensor product spaces.
-In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-order ICT decomposition matrices.
-In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same.
-In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces.
-In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices.
-In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well
 Table 2: Mathematical symbols and their meanings
@@
 $O(n)$ Orthogonal group on $n$-dimensional space $U(n)$ Unitary group on $n$-dimensional space
-$\mathcal{S}$ A vector space $\rho_{\mathcal{S}}$ A group representation
-$\mathcal{S}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{S}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
-$T$ An arbitrary tensor $G\to\mathcal{S}$ $T^{(n)}$ A Cartesian tensor of rank $n$
-$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{S}}$ The set of all possible paths to decompose a tensor in space $\mathcal{S}$
-$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{S};j;q)}$ A general ICT decomposition matrix of space $\mathcal{S}$, weight $j$ and index $q$
+$\mathcal{V}$ A vector space $\rho_{\mathcal{V}}$ A group representation
+$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
+$T$ An arbitrary tensor $G\to\mathcal{V}$ $T^{(n)}$ A Cartesian tensor of rank $n$
+$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
+$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{V};j;q)}$ A general ICT decomposition matrix of space $\mathcal{V}$, weight $j$ and index $q$
 $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
@@
 $P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
-$D^{(\mathcal{S}\to\mathcal{S}^{\prime})}$ A change-of-basis matrix from space $\mathcal{S}$ to $\mathcal{S}^{\prime}$ $v^{(l;p)}$ A vector on which weight $l$ and parity $p$ irreducible representation acts
+$D^{(\mathcal{V}\to\mathcal{V}^{\prime})}$ A change-of-basis matrix from space $\mathcal{V}$ to $\mathcal{V}^{\prime}$ $v^{(l;p)}$ A vector on which weight $l$ and parity $p$ irreducible representation acts
 ${\rm Hom}_{G}$ The set of $G$-equivariant maps from one space to another ${\rm End}_{G}$ The set of $G$-equivariant maps from one space to itself
@@
 $\odot$ Contraction ##\## Number of elements in a set
-as the theory for $SO(n)$, $O(n)$, and $SU(n)$.
+The following contents are organized as follows.
+In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme.
+In Section LABEL:sec:problem, we formulate our problems: obtaining high-rank ICT decomposition matrices and bases for arbitrary tensor product spaces.
+In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-rank ICT decomposition matrices.
+In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same.
+In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces.
+In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices.
+In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$.
 We provide concluding remarks in Section LABEL:sec:conclusion, and introduce the reproducibility details in Section LABEL:sec:repro.
 The acknowledgments are presented after the main text.
-The visualizations are shown in the appendix.
+The complexity comparison with bonvicini2024irreducible is in Appendix LABEL:sec:complexityanalysis.
+The visualizations of ICT decomposition matrices and equivariant bases are shown in Appendix LABEL:sec:visualization.
 A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure LABEL:fig:mindmap.
 2 Preliminaries
+In this section, we recap neccessary mathematical preliminaries.
 2.1 Irreducible Representations and Spherical Spaces
@@
 We first focus on the $O(3)$ group (3-D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion.
-Consider a vector $v\in\mathcal{S}$.
+Consider a vector $v\in\mathcal{V}$.
 The rotation and reflection operations should be able to act on it.
@@
 for any $g_{1}$, $g_{2}\in G$, and * is the group multiplication.
-The group element $g$ acting on the vector $v$ can therefore be represented as a matrix multiplication $\rho(g)\cdot v$.
+The group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$.
 Sometimes, one may also refer to the space of $v$ as a representation space.
-If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{S}$.
-In this sense, this place is also called a $G$-invariant space or subspace.
+If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{V}$, this space is also called a $G$-invariant space or subspace.
 To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal.
@@
 Now, it suffices to define equivariance.
-We claim that a linear function $f:\mathcal{S}\to\mathcal{S}^{\prime}$ is equivariant if and only if
-$\rho_{\mathcal{S}^{\prime}}f(x)=f(\rho_{\mathcal{S}}(x)).$
-We first focus on the simpler case where $\mathcal{S}=\mathcal{S}^{\prime}$.
-We will later generalize this to the case where $\mathcal{S}\neq\mathcal{S}^{\prime}$.
-When it is necessary to emphasize the change in space, we refer to a linear mapping $f$ as $G$-$(\mathcal{S},\mathcal{S}^{\prime})$-equivariant.
+We claim that a linear function $f:\mathcal{V}\to\mathcal{V}^{\prime}$ is equivariant if and only if
+$\rho_{\mathcal{V}^{\prime}}f(x)=f(\rho_{\mathcal{V}}(x)).$
+We first focus on the simpler case where $\mathcal{V}=\mathcal{V}^{\prime}$.
+We will later generalize this to the case where $\mathcal{V}\neq\mathcal{V}^{\prime}$.
+When it is necessary to emphasize the change in space, we refer to a linear mapping $f$ as $G$-$(\mathcal{V},\mathcal{V}^{\prime})$-equivariant.
 We then introduce the definition of reducible and irreducible representation.
 Definition 1
-A group representation $\rho_{\mathcal{S}}:G\to GL(\mathcal{S})$ is said to be reducible, if there is a non-zero subspace $\mathcal{S}^{\prime}\subset\mathcal{S}$, for which we have a representation $\rho_{\mathcal{S}^{\prime}}$, such that for every $g\in G$ and $v\in\mathcal{S}^{\prime}$, it holds that
-$\rho_{\mathcal{S}^{\prime}}(g)v\in\mathcal{S}^{\prime},$
-where we also call $\mathcal{S}^{\prime}$ a $G$-invariant subspace.
-Conversely, if there is no such subspace $\mathcal{S}^{\prime}$ and the corresponding representation on it for a given space $\mathcal{S}$, then we claim that it is irreducible.
+A group representation $\rho_{\mathcal{V}}:G\to GL(\mathcal{V})$ is said to be reducible, if there is a non-zero subspace $\mathcal{V}^{\prime}\subset\mathcal{V}$, for which we have a representation $\rho_{\mathcal{V}^{\prime}}$, such that for every $g\in G$ and $v\in\mathcal{V}^{\prime}$, it holds that
+$\rho_{\mathcal{V}^{\prime}}(g)v\in\mathcal{V}^{\prime},$
+where we also call $\mathcal{V}^{\prime}$ a $G$-invariant subspace.
+Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$, then we claim that $\rho_{\mathcal{V}}$ is irreducible.
 If a representation is irreducible, then the vectors space it acts on is also said to be irreducible. The $3\times 3$ rotation and reflection matrices are irreducible. More generally,

### v7

@@
 The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-rank tensors.
-After decades of research, bonvicini2024irreducible recently achieves an explicit ICT decomposition for $n=5$ with factorial time/space complexity.
-In this work we, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices.
+After decades of research, bonvicini2024irreducible has recently achieved an explicit ICT decomposition for $n=5$ with factorial time/space complexity.
+In this work we, for the first time, obtain decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices.
 The path matrices are obtained via performing chain-like contractions with Clebsch-Gordan matrices following the parentage scheme.
@@
 Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique.
-To the best of our knowledge, this is also the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
+Our method avoids the RREF algorithm and maintains a fully analytical derivation of each ICT decomposition matrix, thereby significantly improving the algorithm’s speed to obtain arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
 We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry.
-The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
+The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-core Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
 Keywords: irreducible Cartesian tensor decomposition, Clebsch-Gordan coefficient, parentage scheme, equivariant graph neural network, equivariant linear mapping
@@
 The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias.
+One typical equivariance requirement is equivariance under the $O(3)$ group (rotations and reflections).
 Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223).
-One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations.
+One of the most promising approaches to constructing equivariant layers involves leveraging irreducible representations.
 Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$.
@@
 where $\hat{T}^{(n;l;q)}$ is an ICT, $\rm vec$} is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $l$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $l$ is sometimes greater than $1$).
+In this paper, we adopt the convention that $\operatorname{vec}$ flattens a tensor by stacking its entries from the rightmost to the leftmost index; concretely, ${\rm vec}(T)=\{T_{...111},T_{...112},\cdots\}$.
 The decomposition matrices $H^{(n;l;q)}$ should be constructed such that $\hat{T}^{(n;l;q)}$ are irreducible.
@@
 However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$.
-Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-rank ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
-It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2).
+Due to the lack of high-rank ICT decomposition matrices, it is impossible to construct high-rank Cartesian EGNNs that incorporate valid equivariant linear layers.
+That is because only irreducible representations of same weight can be linearly combined, but the irreducible representations under Cartesian basis are mixed together.
+In other words, the irreducible representations of the same weight are impossible to be separated, so we cannot linearly combine those of the same type, which limits the available rank of irreducible representations in EGNNs with full access to the equivariant linear layer designs.
+The fact that utilizing high-rank irreducible representations improves performance has been well investigated by predecessors, with the support of several empirical evidence.
+For example, see Table 2 in the paper by batzner20223, where rank $l=3$ NequIP significantly outperforms the same model with lower ranks.
+On the other hand, in the work done by liao2023equiformerv2, Table 1(c) shows that the performance is better with higher ranks.
 Therefore, the matrices for high-rank orthogonal ICT decomposition are of great importance.
-Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
+Based on this, we can obtain valid equivariant linear operations applied to the Cartesian tensors, as these operations constitute the design space of EGNNs.
+In other words, what is the vector space of linear mappings $f$ such that
 $\rho(g)f\left({\rm vec}(T)\right)=f\left(\rho(g){\rm vec}(T)\right)$ (2)
-where $g\in G$, $T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$, and $\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the matrix representation.
+where $g\in G$, an element of a group $G$.
+$T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$ is an arbitrary rank-$n$ tensor.
+$\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the group representation, and $GL(V)$ is a space consisting of the invertible matrices that operate on vector space $V$.
+A detailed introduction to the basic and necessary group theory is in Section 2.1.
 This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$.
@@
 Thus, it is also essential to identify a basis to reduce the number of parameters required.
-In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting.
+In this work, one of our contributions is the proposal of a method that does not rely on RREF algorithm, for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting.
 Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors.
@@
 We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices.
-The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
-If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector.
-Thus, we can multiply it by a diagonal matrix with only zeros and ones on the diagonal.
-This operation eliminates the other spherical vectors and retains only the desired ones.
-The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space.
-The product of these three matrices—the change-of-basis matrix, the diagonal matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs.
-This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz to compute cases for $n=6$, which is already much faster than the current decomposition for $n=5$ by bonvicini2024irreducible.
-We also show that this process can be further accelerated.
-Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that decomposition matrices can be generated via the product of the path matrices and their transpose, with no inverse operation involved.
+The orthogonality of the path matrices further allows us to eliminate the time-consuming inverse operation, rendering this approach an orthogonal decomposition.
+Both theoretical and experimental analyses confirm the efficiency of this approach.
+Inspired by Schur’s lemma, which states that only irreducible representations of the same weight can be linearly combined, we construct an equivariant basis via a path-matrix technique.
+This construction makes it possible to design high-rank equivariant linear layers in EGNNs.
+We then address the lack of a suitable parentage scheme for general spaces.
+To this end, we propose a generalized parentage scheme that yields an equivariant basis for arbitrary spaces.
+These advances collectively enable the efficient construction of equivariant layers across different spaces, thereby expanding the design space of EGNNs.
 Method Rank Orthogonality Time and Space Complexity
@@
 Table 1: Comparison with previous works. The time and space complexities are with respect to rank and dimension, respectively.
-Another challenge this work addresses is the construction of bases for the equivariant spaces.
-With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight).
-These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal.
-Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space.
-We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces.
-We demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities.
-This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
 In summary, this work has five main contributions:
-  • We propose a non-numerical method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
+  • We propose a method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
   • We propose a general parentage scheme to extend ICT decomposition into a generalized ICT decomposition applicable to arbitrary tensor product spaces.
   • We propose to directly obtain orthogonal bases of equivariant spaces, whereas the method proposed by pearce2023brauer yields a spanning set, requiring an expensive post-processing step to extract a basis that is not orthogonal in general.
-  • To the best of our knowledge, this is the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
+  • Our method avoids the RREF algorithm and derives each ICT decomposition matrix in a fully analytical manner, thereby yielding a significantly more efficient algorithm.
   • We extend our approach to efficiently obtain bases of equivariant spaces mapping between different spaces, including spherical tensor product spaces.
@@
 $\mathcal{V}$ A vector space $\rho_{\mathcal{V}}$ A group representation
-$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
+$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;l)}$ The multiplicity of the ICT of rank $n$ and weight $l$
 $T$ An arbitrary tensor $G\to\mathcal{V}$ $T^{(n)}$ A Cartesian tensor of rank $n$
-$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
-$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{V};j;q)}$ A general ICT decomposition matrix of space $\mathcal{V}$, weight $j$ and index $q$
+$\hat{T}^{(n;l;q;p)}$ An ICT of rank $n$, weight $l$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
+[…]
 $C^{l_{1}l_{2}l_{o}}_{m_{1}m_{2}m_{o}}$ A CG coefficient with input weights $l_{1}$, $l_{2}$, output weight $l_{o}$, and the corresponding indices $m_{1}$, $m_{2}$, and $m_{o}$ (the angular-momentum quantum numbers in physics)
 $C^{l_{1}l_{2}l_{o}}$ A CG tensor with input weights $l_{1}$, $l_{2}$, and output weight $l_{o}$ $\hat{C}^{l_{1}l_{2}l_{o}}$ A CG matrix by merging the first two dimensions of the original CG tensor
-$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
+$P^{(path)}$ A path matrix generated by pathpath, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by pathpath
 $D^{(\mathcal{V}\to\mathcal{V}^{\prime})}$ A change-of-basis matrix from space $\mathcal{V}$ to $\mathcal{V}^{\prime}$ $v^{(l;p)}$ A vector on which weight $l$ and parity $p$ irreducible representation acts
@@
 $\otimes$ Tensor product $\oplus$ Direct sum
-$\odot$ Contraction ##\## Number of elements in a set
+$\odot$ Contraction #\# Number of elements in a set
 The following contents are organized as follows.
-In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme.
+In Section 2, we recall basic concepts from representation theory, spherical harmonics, tensors, and tensor product spaces, and we introduce the CG coefficients along with the parentage scheme.
 In Section LABEL:sec:problem, we formulate our problems: obtaining high-rank ICT decomposition matrices and bases for arbitrary tensor product spaces.
@@
 The acknowledgments are presented after the main text.
+The proofs of supportive lemmas are shown in Appendix LABEL:sec:proofs.
 The complexity comparison with bonvicini2024irreducible is in Appendix LABEL:sec:complexityanalysis.
+We give an example for generating rank-2 ICT decomposition matrices in Appendix LABEL:sec:examplerank2.
 The visualizations of ICT decomposition matrices and equivariant bases are shown in Appendix LABEL:sec:visualization.
@@
 2 Preliminaries
-In this section, we recap neccessary mathematical preliminaries.
-2.1 Irreducible Representations and Spherical Spaces
-We first recall some basic facts.
+In this section, we recap necessary mathematical preliminaries.
+2.1 Group, Irreducible Representations and Spherical Spaces
+We first recall some basic facts of group theory.
 A thorough introduction to group theory can be found in inui2012group and brocker2013representations.
-We first focus on the $O(3)$ group (3-D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion.
+A group $G$ is a set equipped with a binary operation *, such that the following five properties hold.
+  • Non-emptiness:
+    The set $G$ is non-empty.
+  • Closure:
+    For all $g_{1},g_{2}\in G$, we have $g_{1}*g_{2}\in G$.
+  • Associativity:
+    For all $g_{1},g_{2},g_{3}\in G$, we have $(g_{1}*g_{2})*g_{3}=g_{1}*(g_{2}*g_{3})$.
+  • Identity element:
+    There exists $e\in G$ such that, for every $g\in G$, it holds that eg=ge=ge*g=g*e=g.
+  • Inverse element:
+    For every $g\in G$ there exists $g^{-1}\in G$ satisfying $g*g^{-1}=g^{-1}*g=e$.
+Group is a powerful tool to study symmetry.
+Rotations and rotations with reflections are both examples of groups.
+For example, it is easy to verify that 2-d rotations satisfy the above requirements: clockwise rotating by $30^{\circ}$, followed by counterclockwise $15^{\circ}$, is equal to clockwise rotating by $15^{\circ}$, and the remained properties can be examined similarly.
+Until now, the group remains abstract, and we cannot bring it into computation.
+Representation theory provides us with a bridge to connect group to linear algebra.
+A representation is a mapping $\rho:G\to GL(V)$ such that
+$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
+holds for any $g_{1},g_{2}\in G$, where $GL(V)$ is a space consisting of the invertible matrices that operate on vector space $V$.
+In other words, they are invertible matrices that operate on vector space $V$.
+Here, we introduce groups that we will consider in this work.
+The orthogonal group, $O(n)$, consisting of orthogonal matrices in $\mathbbm{R}^{n\times n}$, represents $n$-dimensional rotations and reflections; The special orthogonal group, $SO(n)$, consisting of orthogonal matrices with determinant equaling $1$, represents $n$-dimensional rotations; The unitary group, $U(n)$, consisting of unitary complex matrices in $\mathbbm{C}^{n\times n}$, represents $n$-dimensional rotations and reflections on $\mathbbm{C}^{n\times n}$, and $SU(n)$ with additional unit determinant constraint, represents rotations only.
 Consider a vector $v\in\mathcal{V}$.
-The rotation and reflection operations should be able to act on it.
-Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation.
-It holds that
-$\rho(g_{1}*g_{2})=\rho(g_{1})\cdot\rho(g_{2})$
-for any $g_{1}$, $g_{2}\in G$, and * is the group multiplication.
-The group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$.
+A group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$.
 Sometimes, one may also refer to the space of $v$ as a representation space.
 If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{V}$, this space is also called a $G$-invariant space or subspace.
-To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal.
 It is important to note that the co-domain of $\rho$ can differ, meaning that it can also be represented by matrices of other dimensions.
@@
 where we also call $\mathcal{V}^{\prime}$ a $G$-invariant subspace.
-Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$, then we claim that $\rho_{\mathcal{V}}$ is irreducible.
+Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$ satisfying the above conditions, then we claim that $\rho_{\mathcal{V}}$ is irreducible.
 If a representation is irreducible, then the vectors space it acts on is also said to be irreducible. The $3\times 3$ rotation and reflection matrices are irreducible. More generally,